When a NaCl crystal is investigated with x-ray of wavelength 154pm (generated by electron bombardment of a copper target).It was measured that the separation between the neighbouring planes of ion is 2.81×10^-8.Calculate the observed glacing angle
1
Expert's answer
2018-02-14T05:50:26-0500
Wolf-Bragg's formula determines the direction of the diffraction maxima of the X-ray radiation elastically scattered on the crystal:

2dsinƟ = kƛ ,

where d - interplanar distance (lattice period), angle Ɵ is usually called the slip angle, or the angle of reflection, n is the order of the diffraction maximum, λ is the wavelength.

If k=1, sinƟ = ƛ/2d = 154·10-12 m / 2·2.81·10-8 m = 27.4·10-4 = 0.00274

Ɵ = 0.157°

If k=2, sinƟ = ƛ/d = 154·10-12 m / 2.81·10-8 m = 54.8·10-4 = 0.00548

Ɵ = 0.314°